OpenStudy (anonymous):

find the balance in an account with &820 invested at 4% annual simple interest for three years

OpenStudy (anonymous):

I think that it is $922.38. You would get this answer by multiplying the current amount by .04. Then you just do this three times.

OpenStudy (anonymous):

Do you understand?

OpenStudy (triciaal):

Interest I = Principal P x Rate R as a % x Time T in years I = (820)(4/100)(3) = Account balance = starting amount with interest added for this problem 820 + I =